Find the inverse of the function.
f(x)=1+ square root all over x-2.
solve f^-1(x)=(x - blank ^2)+ blank for x greater than equal to blank

Answer by rothauserc(4718)

f(x) = 1 + (x-2)^(1/2)
:
let f(x) = y
:
y = 1 + (x-2)^(1/2)
:
interchange the x's and y's
:
x = 1 + (y-2)^(1/2)
:
solve for y
:
(x-1) = (y-2)^(1/2)
:
square both sides of =
:
(x-1)^2 = y -2
:
y = (x-1)^2 +2
:
substitute f^-1 (x) for y
:
f^(-1) (x) = (x-1)^2 +2 for x > or = 2
: